Accelerator Update for December 21, 2007 - December 31, 2007
Friday December 21
The day shift began with the Tevatron (TeV) in access to investigate a vacuum problem in TeV sector A4. The Antiproton Source (Pbar) stacked, the Recycler stashed, and Operations sent beam to NuMI , SciBooNE / MiniBooNE. The Collider Detector at Fermilab (CDF) and D-Zero (D0) had accessed their collision halls due to the TeV problems. Pbar, Switchyard (SY), and Main Injector (MI) experts conducted studies. MH2 repairs began at 12:30 PM and lasted the rest of the shift. Linac experts began replacing a quadrupole power supply (QPS107) at 1:30 PM and the repairs lasted the rest of the shift.
The evening shift continued with the TeV in access. Experts confirmed that there was a helium leak in the TeV in between the A44 and A45 spool pieces. Vacuum was spoiled and the house warm-up began by mid-shift. Pbar stopped stacking early in the shift and the Accumulator was put into standby mode once it became clear that the TeV would not be ready for beam. Operations transferred the 100 mA stack to Recycler. Linac experts isolated the power supply for QPS107 from its control loop. MI experts conducted MI8 lattice measurements after the Linac was operational. Operators tuned Booster injection and increased efficiency. NuMI and Sci/MiniBooNE took protons when they were available.
Saturday December 22
The midnight shift began with Operations sending beam to NuMI and Sci/MiniBooNE with minimal downtime. The Recycler held on to a stash while the TeV A4 cryo system continued to warm up to room temp. Operators tuned the Booster.
The day and evening shifts had a few 400MeV area trips that operations thought were due to a Linac Klystron cavity (KRF4) LCW valve regulation problem. Operators rebooted a front-end to restore NuMI datalogging. NuMI experts replaced a sump water filter and suffered from the falling barometric pressure that set a number of oxygen sensors alarming.
Sunday December 23
The midnight shift began with NuMI and Sci/MiniBooNE taking beam. KRF4 caused the majority of the Proton Source downtime due to cavity phase temperature regulation problems. The TeV sector A4 cryo system continued to warm up. Operations monitored the Recycler stash. Pbar remained in standby.
The day and evening shifts began with Operations monitoring the Sci/MiniBooNE and NuMI beam, and monitoring the A4 warm up. Operations reported KRF2 and KRF6 trips. Operators tuned Booster and tuned MI with the help of an expert.
Monday December 24
The midnight shift began with Operations sending beam to NuMI and Sci/MiniBooNE despite some Linac power amplifier and Klystron trips. The horizontal positions at the end of the 400 MeV line drifted and required tuning. Pbar remained in standby while the A4 cryo system continued warming.
The day shift began with Operations sending beam to Sci/MiniBooNE and NuMI, with Pbar in standby, and with the Tevatron warming sector A4 for repairs. Experts and Operators prepared TeV sectors A-E for Wednesday's access. Operators turned off NuMI so techs could add water to their LCW system. During this down time, Linac experts replaced a D/A module for QPS107.
The evening shift ended with the TeV A4 cryo system warm and ready for the work to start on Wednesday. Pbar remained in standby mode. The Recycler held its stash. Operations sent beam to the NuMI and Sci/MiniBooNE experiments. KRF4 was off for about two hours. A front-end hard drive failed and experts were called in; repairs continued through the end of shift.
Tuesday December 25
The midnight shift began with Operations sending beam to NuMI and Sci/MiniBooNE. Operators tuned on Booster injection to help decrease losses and increase intensities.
The day shift began with Operations sending beam to NuMI and Sci/MiniBooNE. Operators bypassed BRF1 due to problems, but then brought it back online. A Linac power supply (QPS118) caused problems and KPS1B tripped off a few times. At the end of the shift operators had trouble turning on the MI PS Iinterlock Loop.
The evening shift began with Pbar in standby. Recycler held its stash of 296 E10 antiprotons. Operations sent beam to NuMI and Sci/MiniBooNE. LRF3 and Booster beam position problems caused some downtime. Linac experts worked on LRF3 for about two hours. Operators called a Booster expert for tuning assistance. Operators worked on readback problems for a Linac power supply (QPS118) all shift.
Wednesday December 26
The midnight shift began with Operations sending beam to NuMI and Sci/MiniBooNE. Operators tuned Booster for increased intensities. A TeV expert completed the hipot of sector A4 and dropped the A-E sector interlocks before the A4 work began.
The day shift began with TeV sectors A-E in access for leak checking. Pbar experts conducted stack studies. Recycler experts conducted studies with the stash and then scraped most of it away for further studies. Operations sent beam to NuMI and Sci/MiniBooNE all shift. Operators continued with Linac and Booster tuning and some MI tuning.
The evening shift began with TeV sector A44 mechanical work. Experts verified a liquid helium leak in the vicinity of the A44-5 dipole magnet. Operations sent beam to NuMI and Sci/MiniBooNE. Pbar was in standby mode all shift. Operators tuned Booster and MI for increased efficiency.
Thursday December 27
The midnight shift began with Operations sending beam to Sci/MiniBooNE and NuMI. Operators prepared MI for access, which began at 7 AM. Mechanical techs continued their TeV A44-4 magnet repairs.
The day shift began with TeV A44-4 magnet replacement underway. Vacuum techs accessed MI to replace sections of beampipe. NuMI experts accessed MI-65. The Neutron Therapy Facility (NTF) treated patients. KRF4 tripped off and Linac personnel investigated and had it back on line by 12:30 PM. Linac experts accessed Linac to fix nitrogen leaks on vacuum valves. The Linac access was over by 1:47 PM and then Booster experts established beam to the dump.
The evening shift began with the TeV A44 work continuing and with the MI beam pipe work continuing. Booster experts completed their beam tuning to the dump. Operations reported that the TeV magnet mover was back under the hatch, which meant that Operations could run beam through Transfer Hall when beam was called for. Pbar maintained a stack all shift.
Friday December 28
The midnight shift began with repairs to the 308 section of the MI beam pipe continuing. All of the MI vacuum pumps were on and pumping by 1 AM. Efforts to turn on MI and to established beam were underway, along with sending beam to Sci/MiniBooNE and NuMI, by 2 AM. Operators changed the Klystron 7 phase, which corrected Booster beam position problems and significantly increased efficient delivery of beam to MI. A power supply (HV703) developed an LCW leak and delayed beam to Pbar.
The day shift began with work continuing on the TeV A44-4 dipole replacement. Vacuum experts continued their MI access to search for more vacuum leaks. Pbar techs and experts continued with HV703 repairs and a lithium lens investigation. The Pbar AP0 lens change out began around 10:30 AM. Operations put the TeV A0 kickers into condition mode. TeV experts hipotted sector A. The MI access ended at 1 PM and beam was restored to NuMI and MiniBooNE by 1:45 PM. Experts replaced all the TeV A0 kicker filaments. Operations sent beam to the Recycler. HV703 repairs lasted all shift.
The evening shift began with the continuing TeV work. Pbar experts completed the lens replacement around 9 PM and about 30 minutes later began a low intensity stacking target tune up. Operations sent beam to NuMI and Sci/MiniBooNE when it was available.
Saturday December 29
The midnight shift began with the TeV A44 vacuum repairs continuing. Operators tuned Booster, MI, and Pbar for increased intensity and better beam positions. Antiprotons were stacked and transferred four times to the Recycler, after experts setup the target station. Operations sent beam to the NuMI and Sci/MiniBooNE experiments with minimal downtime.
The day and evening shifts continued with TeV sector A44-4 repairs, stacking, and with beam to NuMI and Sci/MiniBooNE. Mechanical Support experts completed their work and turned over A44 to cryo system experts to reconnect the magnet string to the cryo system. This work began at 2:30 PM and ended at 7 PM. Pbar experts conducted studies and the Recycler stashed once during the shift. KRF6 tripped off at the end of shift.
Sunday December 30
The midnight shift began with the completion of TeV A4 mechanical sign-offs. The cryo cool down, which should take about 24 hours, started by mid-shift. Operators searched and secured the Tevatron. Operations stacked and delivered beam to the NuMI and Sci/MiniBooNE experiments all night with minimal downtime.
The day and evening shifts began with the TeV A4 cryo system in cooldown mode, with Pbar stacking, with Recycler stashing, and with beam to the NuMI and Sci/MiniBooNE experiments. TeV A4 signoffs occurred in the morning and permission to power the TeV bus was given by 11 AM. A problem with the B0Q3 HFU discharge times was due to the temperature of the magnet. The HFU test was successfully completed later in the shift. Operators encountered two Pbar stacking problems. VT714 was replaced but did not put out the requisite current until it was power cycled. The Accumulator unstacking was fixed with a -6 Hz change to the stabilizing frequency. Experts accessed the TeV from 3 to 4 PM to stop the E0 bake and remove the magnet mover at D0. A D0 collision hall access began at 6 PM and lasted the rest of the shift.
Monday December 31
The midnight shift began with TeV sector A4 cryo system going cold by 1:30 AM. Operators prepared the TeV to take beam. There were various problems with the A3 FB Loop input and the TeV CAMAC controls that slowed turn-on. Operations had TeV beam to flattop by the end of shift. The Accumulator held its stack in standby mode and the Recycler held its stash. Operations sent beam to the NuMI and Sci/MiniBooNE experiments when beam was available. The Sci/MiniBooNE horn tripped once with a ground fault indication; it reset without incident. The Linac I- Source vacuum began to degrade toward the end of shift. Various Linac Klystrons tripped with CS fault indications. Booster experts tuned for increased efficiency at the end of shift.
The day shift began with TeV studies, stacking, stashing, beam to the NuMI and Sci/MiniBooNE experiments. CDF made an access into their collision hall. TeV studies continued once the access was over. The Recycler stashed twice. Operations put the TeV into shot setup at 1:44 PM and established store 5819 at 3:22 PM with an initial luminosity of 193.22e30.
During the evening shift the Recycler stashed four times. Pbar had problems with its autotune program, which took a while to work out. Operations sent beam to NuMI and Sci/MiniBooNE when it was available. TeV store 5819 collided all shift.
The Shift Summary for the New Year through to January 6 th will be on Wednesday
Comments and Suggestions
|last modified 01/08/2008 email Fermilab|
|Security, Privacy, Legal|